Two brothers have the following physical quantities ( 400an, 2.5m) and (600N, 2.0m) respectively in a walking exercise. They both exerted a total kinetic energy of 20,000Joules in a 100m race competition with the second brother possessing a kinetic energy of 7500joules.

Use the information to answer the following questions:

1. What was the amount of work done by both of them?

2. What was the value of velocity or speed of the first brother during the race

3. Determine the size of the masses when they both walked.

4. State the summation value of their potential energy in kilojoules
1
Expert's answer
2020-07-06T15:12:23-0400

1)


"W=F_1s+F_2s\\W=(400)(100)+(600)(100)=100000\ J\\W=100\ kJ"

2)

"0.5(2m)v_2^2=7500\\0.5(2m)v_2^2+0.5(2.5m)v_1^2=20000"

"7500+0.5(2.5m)v_1^2=20000"

"v_1=22.4\frac{m}{s}"

3)


"m_1=2.5(20)=50\ kg\\m_2=2(20)=40\ kg"

4)


"U=W-K=100-20=80\ kJ"